# Non-isothermal diffuse interface model for phase transition and interface evolution

### v1

@@ 4 Numerics
In this section, we perform a numerical study of the non-isothermal Allen-Cahn equation (3.1), which can be rewritten as
(4.1) $\begin{cases}&\varphi_{t}=-\frac{1}{\epsilon}\mu,\quad\mu=\frac{1}{\epsilon}(\varphi^{2}-1)\varphi-\epsilon\Delta\varphi+\frac{\chi}{\theta_{c}}(\theta-\theta_{c})(\varphi^{2}-1).\\ &c_{v}\theta_{t}=\nabla\cdot(\kappa(\varphi)\nabla\theta)-\frac{\chi\gamma}{\theta_{c}}\theta(\varphi^{2}-1)\mu+\gamma|\mu|^{2}\\ \end{cases}$
The numerical simulations focus on the effects of the curvature in the dynamics, with $\chi=1$. Our aim is to validate the results of the asymptotic expansion discussed in the previous section.
@@
(4.2) $c_{v}\theta_{0}\tilde{\theta}_{t}=\theta_{0}\nabla\cdot(\kappa\nabla\tilde{\theta})-\frac{\chi\gamma}{\tilde{\theta}_{c}}\tilde{\theta}(\varphi^{2}-1)\mu+\gamma|\mu|^{2}\ ,\$
@@
To solve (4.1), we adopt the following semi-implicit scheme for temporal discretization
  • Fixed $\theta^{n}$, solve equation of $\varphi$ using the implicit Euler scheme
    (4.3) $\frac{\varphi^{n+1}-\varphi^{n}}{\Delta t}=-\gamma\mu(\varphi^{n+1},\theta^{n})\ .$
  • With $\varphi^{n+1}$, solve the heat equation using
    (4.4) $c_{v}\frac{\theta^{n+1}-\theta^{n}}{\Delta t}=\nabla\cdot(\kappa\nabla\theta^{n+1})-\frac{\chi\gamma}{\theta_{c}}\mu^{n+1}g^{\prime}(\varphi^{n+1})\theta^{n+1}+\gamma|\mu^{n+1}|^{2}\ ,$
@@
    (4.5) $\mu^{n+1}=\frac{1}{\epsilon}((\varphi^{n+1})^{2}-1)\varphi^{n+1}-\epsilon\Delta\varphi^{n+1}+\frac{\chi}{\theta_{c}}(\theta^{n}-\theta_{c})((\varphi^{n+1})^{2}-1)\ .$
We use the standard central finite difference scheme for the spatial discretization. The current numerical scheme is quite simple and may not preserve the original energetic variational structure in the discrete level. We plan on developing a structure-preserving numerical scheme to (4.1) in the future work.
In all numerical simulations below, we take $\epsilon=0.05$, $h=1/50$ and $\Delta t=10^{-3}$.
@@ 4.1 Quasi-1D case
We first consider a quasi-1D case, by taking the initial condition as $\varphi_{0}(x,y)=\tanh\left(\frac{x+0.1\sin(10y)}{\sqrt{2}\epsilon}\right)$ and $\theta_{0}=\theta_{c}+10\varphi_{0}(x,y)$. We impose the Neumann boundary condition for both $\varphi$ and $\theta$. Fig. 1 shows the numerical results for different times $t$. One can notice that the movement of the interface $\theta=\theta_{c}$ is almost same to the movement of the interface $\varphi=0$, both are driven by the curvature effect.
The initial condition is motivated by the numerical example in [16], which can be used to explore the Mullins-Sekerka instability in some Stefan problems. It is worth mentioning that the Allen-Cahn type phase field model mainly captures the curvature effects in the interfacial dynamics. To capture the Mullins-Sekerka instability in some Stefan problems, one should study a non-isothermal Cahn-Hilliard type phase-field model [23].
@@ 4.2 Motion by mean curvature
Next, we consider the initial conditions $\varphi_{0}({\bm{x}})=\tanh\left(\frac{\sqrt{x^{2}+y^{2}-0.6}}{\sqrt{2}\epsilon}\right)$ and $\theta_{0}({\bm{x}})=10\varphi_{0}({\bm{x}})+273.15$, where we impose Dirichlet boundary conditions to both $\varphi$ and $\theta$.
Fig. 2(a)-(d) shows the numerical results in this case with $\epsilon=0.05$ at different time. Fig. 2(e) shows the movement of interfaces $\varphi=0$ and $\theta=\theta_{c}$. It is well known that in the isothermal case, the dynamics of the interface in the phase-field model will converge to the mean curvature flow, governed by $R(t)=\sqrt{0.36-2t}$, as $\epsilon\rightarrow 0$. Here, $R(t)$ represents the radius of the interface. The numerical results clearly suggest that the movement of interfaces $\varphi=0$ and $\theta=\theta_{c}$ follows the mean curvature flow when the time $t$ is small, which is consistent with the asymptotic analysis.
Figure 2: Numerical results for $\epsilon=0.05$ at different time (left: $\varphi$ ; right: $\theta$ ): (a) $t=0$, (b) $t=0.01$, (c) $t=0.05$, and (d) $t=0.1$. (e) Location of interface $\varphi=0$ and $\theta=\theta_{c}$ as a function of time.
Next, we consider more complicated initial shapes for the droplets. Fig. 3 shows the numerical results for the initial condition $\varphi_{0}({\bm{x}})=\tanh\left(\frac{\sqrt{x^{2}+4y^{2}-0.6}}{\sqrt{2}\epsilon}\right)$ and $\theta_{0}({\bm{x}})=10\varphi_{0}({\bm{x}})+273.15$, along with Dirichlet boundary conditions. Again, the movement of both interfaces are driven by the mean curve, and almost same initially. But the interface $\theta=\theta_{c}$ moves faster and disappears early.
Figure 3: Numerical results for $\epsilon=0.05$ at different time [left: $\varphi(x,t)$, right: $\theta$]: (a) $t=0$, (b) $t=0.01$, (c) $t=0.02$, (d) $t=0.03$, (e) $t=0.05$, and (f) $t=0.07$.
Figure 4: Numerical results for $\epsilon=0.05$ at different time [left: $\varphi({\bm{x}},t)$, right: $\theta$]: (a) $t=0$, (b) $t=0.01$, (c) $t=0.02$, (d) $t=0.05$, (e) $t=0.07$, and (f) $t=0.1$.
@@
We also consider an initial condition with a triangular shape of the droplet, shown in Fig. 4(a). The simulation results are shown in Fig. 4(a) - (f) for different times $t$. Similar to the previous case, the movement of the two interfaces is in agreement when $t$ is small, but will diverge from each other for larger times.
@@ 4.3 Beyond the Stefan problem
As mentioned earlier, in order to reduce the non-isothermal Allen-Cahn equation to the classical Stefan problem in the sharp interface limit, it is crucial to choose the proper scaling of $\chi(\epsilon)$ and $\gamma(\epsilon)$ in the asymptotic analysis. Fig. 5 shows a simulation result for $\chi(\epsilon)=\frac{1}{\epsilon}$. The initial and boundary conditions are same to that in Fig. 2. Under this scaling, the movement of interfaces of the temperature field and phase function are no longer consistent with each other even in short time. Although the curvature effects still dominate the motion of the interface of $\varphi$, the movement of the interface is slower than the mean curvature flow. In contrast to the Stefan problem, the movement of interface of the temperature no longer follows the mean curvature flow.
Figure 5: Numerical results for $\epsilon=0.05$ at different time (left: $\varphi$ ; right: $\theta$ ): (a) $t=0$, (b) $t=0.01$, (c) $t=0.05$, and (d) $t=0.1$. (e) Location of interface $\varphi=0$ and $\theta=\theta_{c}$ as a function of time.
@@ 5 Conclusion
In this paper, we derive a non-isothermal Allen-Cahn model for phase transition and heat transfer by using an energetic variational approach. In contrast to the Stefan problem, which employs a single interface to describe phase transition and temperature, the non-isothermal Allen-Cahn model contains more physical information. By asymptotic analysis, we have shown that the sharp interface limit of the non-isothermal Allen-Cahn equation, under certain scaling of the melting/freezing energy, results in a two-phase Stefan like problem. Various numerical simulations further demonstrate that the Stefan problem is a good approximation to the non-isothermal Allen-Cahn equation in the short time, as evidenced by the close agreement in the evolution of two interfaces. However, over longer time periods, the movement of two interfaces will diverge from each other.
While diffuse interface models can be used to study different effects of interface dynamics during phase transitions, the Allen-Cahn type model examined in this paper mainly focuses on curvature effects. In contrast, other physical phenomena, such as the Mullins-Sekerka instability, can only be captured by the Cahn-Hilliard type diffuse interface model [23]. We will perform numerical studies and analyze the asymptotic limit of these diffuse interface models in future work.
@@ Acknowledgements
C. L. was partially supported by NSF DMS-2153029 and DMS-2118181. J.-E. S. was supported by the German Science Foundation (DFG) under grant No.456754695. Y. W. was supported by NSF DMS-2153029.

### v2

@@ 4 Numerics
In this section, we perform a numerical study of the non-isothermal Allen-Cahn equation (3.1) - (3.2) with $\lambda=1$, which can be rewritten as
(4.1) $\begin{cases}&\phi_{t}=-\frac{1}{\epsilon}\mu,\quad\mu=\frac{1}{\epsilon}(\phi^{2}-1)\phi-\epsilon\Delta\phi+\frac{1}{\theta_{c}}(\theta-\theta_{c})(\phi^{2}-1).\\ &c_{v}\theta_{t}=\nabla\cdot(\kappa(\phi)\nabla\theta)-\frac{1}{\epsilon}\frac{\theta}{\theta_{c}}(\phi^{2}-1)\mu+\frac{1}{\epsilon}|\mu|^{2}\\ \end{cases}$
The numerical simulations focus on the effects of the curvature in the dynamics. Our aim is to validate the results of the asymptotic expansion discussed in the previous section and provide the behavior of solution for finite $\epsilon$.
@@
(4.2) $c_{v}\theta_{0}\tilde{\theta}_{t}=\theta_{0}\nabla\cdot(\kappa\nabla\tilde{\theta})-\frac{1}{\tilde{\theta}_{c}\epsilon}\tilde{\theta}(\phi^{2}-1)\mu+\frac{1}{\epsilon}|\mu|^{2}\ ,$
@@
To solve the equation (4.1), we adopt the following semi-implicit scheme for temporal discretization:
  • Fixed $\theta^{n}$, solve equation of $\phi$ using the implicit Euler scheme
    (4.3) $\frac{\phi^{n+1}-\phi^{n}}{\Delta t}=-\frac{1}{\epsilon}\left(\frac{1}{\epsilon}((\phi^{n+1})^{2}-1)\phi^{n+1}-\epsilon\Delta\phi^{n+1}+\frac{1}{\theta_{c}}(\theta^{n}-\theta_{c})((\phi^{n+1})^{2}-1)\right)$
  • With $\phi^{n+1}$, solve the heat equation using
    (4.4) $c_{v}\frac{\theta^{n+1}-\theta^{n}}{\Delta t}=\nabla\cdot(\kappa\nabla\theta^{n+1})-\frac{1}{\theta_{c}\epsilon}\mu^{n+1}g^{\prime}(\phi^{n+1})\theta^{n+1}+\frac{1}{\epsilon}|\mu^{n+1}|^{2}\ ,$
@@
    $\mu^{n+1}=\frac{1}{\epsilon}((\phi^{n+1})^{2}-1)\phi^{n+1}-\epsilon\Delta\phi^{n+1}+\frac{1}{\theta_{c}}(\theta^{n}-\theta_{c})((\phi^{n+1})^{2}-1).$
We use the standard central finite difference scheme for the spatial discretization. The current numerical scheme is quite simple and may not preserve the original energetic variational structure in the discrete level. We plan on developing a structure-preserving numerical scheme to (4.1) in the future work. In all numerical simulations below, we take $\epsilon=0.05$, $h=1/50$ and $\Delta t=10^{-3}$ unless stated otherwise.
@@ 4.1 Quasi-1D case
We first consider a quasi-1D case, by taking the initial condition as
$\phi_{0}(x,y)=\tanh\left(\frac{x+0.1\sin(10y)}{\sqrt{2}\epsilon}\right),\quad\theta_{0}=\theta_{c}+10\phi_{0}(x,y)$
We impose the Neumann boundary condition for both $\phi$ and $\theta$. Fig. 1 shows the numerical results for different times $t$. One can notice that the movement of the interface $\theta=\theta_{c}$ is almost same to the movement of the interface $\phi=0$, both are driven by the curvature effect.
The initial condition is motivated by the numerical example in [16], which can be used to explore the Mullins-Sekerka instability in some Stefan problems. It is worth mentioning that the Allen-Cahn type phase field model mainly captures the curvature effects in the interfacial dynamics. To investigate the Mullins-Sekerka instability and dendritic growth, one should study a non-isothermal Cahn-Hilliard type phase-field model [23].
@@ 4.2 Motion by mean curvature
Next, we consider the initial conditions
(4.5) $\phi_{0}({\bm{x}})=\tanh\left(\frac{\sqrt{x^{2}+y^{2}-0.6}}{\sqrt{2}\,\epsilon}\right),\qquad\theta_{0}({\bm{x}})=10\,\phi_{0}({\bm{x}})+273.15,$
where Dirichlet boundary conditions are imposed on both $\phi$ and $\theta$.
A Initial condition (4.5)
B Initial condition (4.6)
[…]
Figure 2A(a)–(d) displays the numerical results for $\epsilon=0.05$ at different times, while Figure 2A(e) shows the evolution of the interfaces defined by $\phi=0$ and $\theta=\theta_{c}$. The asymptotic analysis presented earlier shows that, in the sharp-interface limit, the motion of both interfaces converges to mean curvature flow, described by
$R(t)=\sqrt{0.36-2t},$
as $\epsilon\to 0$, where $R(t)$ denotes the radius of the interface. Numerical results confirm that, for small times $t$, the interfaces $\phi=0$ and $\theta=\theta_{c}$ both evolve according to mean curvature flow, in agreement with the asymptotic prediction. For larger times, however, the two interfaces gradually diverge, exhibiting a superheating effect in which the temperature within certain regions of the solid phase exceeds the phase-transition threshold. We also consider the initial condition
(4.6) $\phi_{0}({\bm{x}})=-\tanh\left(\frac{\sqrt{x^{2}+y^{2}-0.6}}{\sqrt{2}\,\epsilon}\right),\qquad\theta_{0}({\bm{x}})=10\,\phi_{0}({\bm{x}})+273.15,$
with results shown in Fig. 2B. Similarly, the $\phi$-interface follows mean curvature flow, while the $\theta=\theta_{c}$ interface deviates from it. Superheating is again observed, as the temperature in certain regions of the solid phase remains above the critical temperature.
To further demonstrate the numerical convergence of the phase-field model to its sharp-interface limit, we consider four different $\epsilon$: $\epsilon=0.1,0.05,0.02$ and 0.010.010.01. Figures 3(a)-(b) show the evolution of interfaces for $\phi=0$ and $\theta=\theta_{c}$ for different $\epsilon$, respectively. As expected, the $\phi=0$ interface approaches the mean curvature flow as $\epsilon\to 0$, which is consistent with the asymptotic analysis. The evolution of the interface $\theta=\theta_{c}$ is in agreement with mean curvature flow in the short time, but deviate from $\phi=0$ even for small $\epsilon$ when $t$ goes large, leading to the superheating phenomenon. The gap between the two interfaces does not visibly shrink as $\epsilon$ decreases.
[…]
Figure 3: Evolution of the interfaces: (a) $\phi=0$ and (b) $\theta=\theta_{c}$, for varying $\epsilon$.
The numerical result can be understood as follows: the phase-field equation at $\mathcal{O}(1/\epsilon)$ enforces the Dirichlet condition $\theta=\theta_{c}$ on the moving interface. In principle, as $\epsilon\to 0$, the two interfaces are expected to coincide. Nevertheless, the evolution of the temperature field is governed by a heat equation at $\mathcal{O}(1)$, which adjusts only diffusively in time and is not strictly slaved to the motion of the $\phi$-interface. Consequently, the $\theta=\theta_{c}$ level set may propagate faster than the $\phi=0$ interface, leading to the superheating effect. This scale separation between the phase-field dynamics and the heat equation likely explains why the numerical simulations do not exhibit a clear collapse of the gap between the two interfaces, even for relatively small $\epsilon$. In practice, resolving this discrepancy would require significantly smaller values of $\epsilon$, which are beyond the range currently tractable by the numerical methods employed.
Next, we consider more complicated initial shapes for the droplets. Fig. 4A shows the numerical results for the initial condition
(4.7) $\phi_{0}({\bm{x}})=\tanh\left(\frac{\sqrt{x^{2}+4y^{2}-0.6}}{\sqrt{2}\epsilon}\right),\quad\theta_{0}({\bm{x}})=10\phi_{0}({\bm{x}})+273.15\ ,$
@@
along with Dirichlet boundary conditions. Again, the movement of both interfaces are driven by the mean curve, and almost same initially. But the interface $\theta=\theta_{c}$ moves faster and disappears early. The evolution at different snap shots in time is shown in Fig 4A (a)-(f). We also consider an initial condition with a triangular shape of the droplet, shown in Fig. 4B. The simulation results are shown in Fig. 4B(a) - (f) for different times $t$. Similar to the previous case, the movement of the two interfaces is in agreement when $t$ is small, but will diverge from each other for larger times.
A Evolution of initial condition (4.7)
B Evolution of triangular initial condition
[…]
@@ 4.3 Beyond the Stefan problem
As mentioned earlier, in order to reduce the non-isothermal Allen-Cahn equation to the classical Stefan problem in the sharp interface limit, it is crucial to choose the proper scaling of $\chi(\epsilon)$ in the asymptotic analysis. Fig. 5 shows a simulation result for $\chi(\epsilon)=\frac{1}{\epsilon}$. The initial and boundary conditions are same to that in Fig. 2A. Under this scaling, the movement of interfaces of the temperature field and phase function are no longer consistent with each other even in short time. Although the curvature effects still dominate the motion of the interface of $\phi$, the movement of the interface is slower than the mean curvature flow. In contrast to the Stefan problem, the movement of interface of the temperature no longer follows the mean curvature flow.
Figure 5: Numerical results for $\epsilon=0.05$ at different time (left: $\phi$ ; right: $\theta$ ): (a) $t=0$, (b) $t=0.01$, (c) $t=0.05$, and (d) $t=0.1$. (e) Location of interface $\phi=0$ and $\theta=\theta_{c}$ as a function of time.
@@ 5 Conclusion
In this paper, we derive a non-isothermal Allen-Cahn model for phase transition and heat transfer by using an energetic variational approach. In contrast to the Stefan problem, which employs a single interface to describe phase transition and temperature, the non-isothermal Allen-Cahn model contains more physical information. By asymptotic analysis, we have shown that the sharp interface limit of the non-isothermal Allen-Cahn equation, under certain scaling of the melting/freezing energy, results in a nonlinear Stefan type problem. Various numerical simulations further demonstrate that the Stefan problem is a good approximation to the non-isothermal Allen-Cahn equation in the short time, as evidenced by the close agreement in the evolution of two interfaces. However, over longer time periods, the movement of two interfaces will diverge from each other.
While diffuse interface models can be used to study different effects of interface dynamics during phase transitions, the Allen-Cahn type model examined in this paper mainly focuses on curvature effects. In contrast, other physical phenomena, other physical phenomena, such as the Mullins-Sekerka instability and dendritic growth, can only be captured by the Cahn-Hilliard type diffuse interface model [23]. We will perform numerical studies and analyze the asymptotic limit of these diffuse interface models in future work.
@@ Acknowledgements
C. L. was partially supported by NSF DMS-2153029, DMS-2118181 and DMS-2410742. J.-E. S. was supported by the German Science Foundation (DFG) under grant No.456754695. Y. W. was supported by NSF DMS-2153029 and DMS-2410740.